OpenStudy (anonymous):

-21=-3/5t find t (this is a fraction but i don't know how to insert it so i put the slash sign :)

OpenStudy (anonymous):

-3/( 5t ) ?

OpenStudy (anonymous):

is it 3/(5t) or (3/5)t?

OpenStudy (anonymous):

If \[\frac{ 3}{ 5t}\] then answer is 1/35 If (3/5) t then answer is 35

OpenStudy (anonymous):

\[-21=\frac{ -3}{ 5}t\]

OpenStudy (anonymous):

this is the equation

OpenStudy (anonymous):

how do answer that, step by step?

OpenStudy (anonymous):

i have to find t

OpenStudy (anonymous):

all you need to do is isolate the t... clear the fraction on the right side by multiplying by the reciprocal of -3/5 (which is -5/3): \(\LARGE -21=\frac{-3}{5}t \) \(\LARGE \color {red}{\frac{-5}{3}} \cdot -21=\color {red}{\frac{-5}{3}} \cdot \frac{-3}{5}t \) can you take it from here?

OpenStudy (anonymous):

uhmm can u do the rest im a bit confused..

OpenStudy (anonymous):

i have to switch thefraction so i can mulktiply?

OpenStudy (anonymous):

fraction so i can multiply?

OpenStudy (anonymous):

t= -35 ???

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

notice on the left side you multiplied a negative number by another negative number... your left side should be positive...

OpenStudy (anonymous):

positive 35....

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

dude ur the best at expalining pls be my personal tutor person on here xD